1.If a worker is exposed to 2 hours at 85 dBA, 1 hour at 90 dBA, 4 hours at 88 dBA and 1 hour at 100 dBA, has the OSHA noise exposure standard been exceeded?

2. If a worker has been exposed for 2 hours at 95 dBA at the beginning of the workshift, for how many hours can the person be exposed to 100 dBA that same day?
Answer #1

1)
we know, OSHA's permissible exposure limit (PEL) is 90 dBA for all workers for an 8 hour day.

In one day maximum noice exposure = 90*8

= 720 dBA.hour

In the given case

noise exposure = 2*85 + 1*90 + 4*88 + 1*100

= 712 dBA.hour

clearly, 712 < 720

so, OSHA noise exposure standard not been exceeded.

2) let t is the required hours.

720 = 95*2 + 100*t

==> t = (720 - 95*2)/100

= 5.3 hours